# Hybrid-Emba3D: Geometry-Aware and Cross-Path Feature Hybrid Enhanced State Space Model for Point Cloud Classification

### v1

@@ Abstract
The point cloud classification tasks face the dual challenge of efficiently extracting local geometric features while maintaining model complexity. The Mamba architecture utilizes the linear complexity advantage of state space models (SSMs) to overcome the computational bottleneck of Transformers while balancing global modeling capabilities. However, the inherent contradiction between its unidirectional dependency and the unordered nature of point clouds impedes modeling spatial correlation in local neighborhoods, thus constraining geometric feature extraction. This paper proposes Hybrid-Emba3D, a bidirectional Mamba model enhanced by geometry-feature coupling and cross-path feature hybridization. The Local geometric pooling with geometry-feature coupling mechanism significantly enhances local feature discriminative power via coordinated propagation and dynamic aggregation of geometric information between local center points and their neighborhoods, without introducing additional parameters. The designed Collaborative feature enhancer adopts dual-path hybridization, effectively handling local mutations and sparse key signals, breaking through the limitations of traditional SSM long-range modeling. Experimental results demonstrate that the proposed model achieves a new SOTA classification accuracy of 95.99% on ModelNet40 with only 0.03M additional. Our code and weights are available at https://github.com/L1277471578/Hybrid-Emba3D.
@@
State Space Model, Point Cloud, Classification, Local geometry
@@
[label1] organization=Xi’an Technological University,city=Xi’an, postcode=710021, state=Shaanxi, country=China
@@
[label2]organization=Changchun University of Science and Technology, city=Changchun, postcode=130022, state=Jilin, country=China
@@ 1 Introduction
Point clouds directly represent the geometric features of objects through 3D coordinates, compensating for the lack of spatial information in 2D data and providing basic data support for perception tasks such as 3D object recognition and scene understanding[19]. However, the inherently disordered nature of point cloud data necessitates permutation invariant processing, and effective extraction of local geometric features is constrained by nonuniform distribution, these intrinsic characteristics present challenges for improving the accuracy and generalization capability of 3D point cloud classification models.
In response to the challenges of point cloud disorder and local feature extraction, researchers have explored. Among them, PointNet[4] innovatively integrates multi-layer perceptrons and symmetric functions, and builds a processing framework with permutation invariance for the first time, laying the foundation for the direct representation of the disordered features of point clouds. Subsequent research further captures local geometric features by developing point cloud convolution operations[44, 20, 47]. However, the convolution-based method is essentially local feature aggregation, which makes it difficult to model the long-distance dependency in point cloud data and limits its performance improvement in complex scenes. Some researchers have shifted their focus to Transformer[38] architectures with global modeling capabilities to overcome the limitations of the convolution method.
The PointTransformer[8] pioneered the application of self-attention mechanisms in 3D point cloud processing, utilizing positional encoding and multi-head attention mechanisms to establish comprehensive modeling of global dependencies within point clouds. While subsequent studies have enhanced feature representation capabilities through refined local regional modeling and hierarchical architectural designs, the intrinsic quadratic computational complexity of Transformer-based frameworks persistsently suffers from dual constraints on computational resources and inference speed[18].
@@
State Space Model (SSM) is an emerging sequence modeling method in the field of deep learning. It realizes efficient long sequence modeling with the mathematical characteristics of continuous state representation and linear time-invariant systems. By introducing a selective mechanism with input dependency, the Mamba model[11] innovatively constructs a state space architecture with conditional computing characteristics. It combines the timing dependence modeling capabilities of RNNs[22] with the parallel computing advantages of the Transformer to achieve linear time complexity of long-distance dependence modeling. However, the irregular topological characteristics of point cloud data make it difficult for the SSM discrete state transfer mechanism to effectively capture geometric correlations. Furthermore, SSM’s unidirectional scanning paradigm is inadequate in modeling the intrinsically bidirectional spatial dependencies embedded within point clouds, particularly in local feature refinement processes where this limitation inevitably induces structural information loss of geometric details.
To address these limitations, this paper proposes a Geometry-Aware and Cross-Path Feature Hybrid Enhanced Mamba Model (Hybrid-Emba3D), which realizes lightweight local spatial information feature enhancement and feature cross-path fusion, improving the ability to extract local and global information.
@@
    A lightweight feature coupling mechanism based on geometric priors is proposed, which couples geometric feature information between center point features and neighboring point spatial geometric information, enriching local geometric representations.
@@
    Designing a multi-dimensional hybrid feature enhancement module with channel-wise and multi-scale learning capabilities, which enhances context-aware modeling capacity for SSM.
@@
    The proposed model demonstrates exceptional performance, particularly outperforming existing SOTA models on the ModelNet40 benchmark and establishing new performance benchmarks. As shown in Figure 1.
Figure 1: Comprehensive comparison of Hybrid-Emba3D with other SOTA models
@@ 2.1 Deep learning for point cloud data
Effective point cloud classification requires the collaborative modeling of local geometric details and global spatial context. Maturana[29] et al. proposed VoxNet, which organizes unstructured point clouds into 3D voxel grids and applies conventional 3D convolutional neural networks to capture comprehensive spatial contextual features. Based on PointNet, PointNet++[31] employs Farthest Point Sampling (FPS) to construct hierarchical architectures while enhancing localized detail extraction through its local feature aggregation mechanism. Its density-adaptive modules effectively mitigate multi-scale learning interference caused by non-uniform point density distributions. Subsequent approaches encompass point-based direct convolution[36, 26, 9, 42], Volume-Based convolution [34, 35, 17, 43], and graph-based convolution[41, 28, 40, 24]. They are all attempts to enhance local feature extraction through spatial transformations or deformable convolutional kernels. However, these methodologies fundamentally suffer from the mismatch between fixed receptive fields and irregular point distributions.
In 2021, PointTransformer[8] was the first to introduce the Transformer architecture, marking a new stage in global context modeling for point cloud processing. Based on relative position coding, its self-attention mechanism can adaptively establish geometric correlations between point cloud elements and significantly improve the scene semantic understanding ability through global dependency modeling[30, 48, 51, 14, 45]. Notably, the architecture’s intrinsic permutation equivariance naturally conforms with the unordered nature of point clouds, eliminating the need for data rearrangement or interpolation required by voxel/grid-based methods, thereby avoiding the information loss caused by three-dimensional coordinate discretization.
@@
The conflict exists between the sparse distribution characteristics of point cloud data and the dense connectivity assumption underlying standard attention mechanisms, leading to redundant attention weight allocations. Researchers have developed sparse attention strategies to improve this. Such as DSVT[10] with hierarchical feature aggregation demonstrates optimized computation-accuracy balance through adaptive sparse attention patterns; Point-BERT[48] self-supervised pretraining via masked point modeling, effectively reducing annotation dependency; PatchFormer[50] significantly reduces computational complexity through geometrically consistent patch partitioning and streamlined patch-level attention mechanisms.
While convolution-based architectures exhibit linear computational complexity, their localized receptive fields fundamentally constrain long-range contextual modeling capabilities. Conversely, Transformer models achieve global interaction through self-attention mechanisms, yet are inherently limited by quadratic computational complexity. In this context, state space models (SSMs) like Mamba demonstrate potential for point cloud processing, achieving linear modeling characteristics in long-sequence modeling that significantly reduce both computational and spatial overhead while preserving global dependency learning.
2.2 State Space Model
@@
SSM, a classic method in control theory, achieves dynamic system modeling by constructing the state space. As control theory evolved, the limitations of traditional input-output models have gradually emerged. SSM establishes a system state space based on input signal characteristic variables, uses state variables to characterize the dynamic evolution of the system fully, and abstracts the system into a mathematical model containing state and output equations. The SSM structure are shown in Figure 2.
@@
Discretize the continuous-time linear time-invariant (LTI) state space model using the zero-order hold (ZOH) method. The discretized equation of state is given in equation 1.
${}\left\{\begin{matrix}{{h}_{k}}&=\bar{A}\cdot{{h}_{k-1}}+\bar{B}\cdot{{x}_{k}}\\ {{y}_{k}}&=\bar{C}\cdot{{h}_{k}}+\bar{D}\cdot{{x}_{k}}\\ \end{matrix}\right.$ (1)
Where: $\bar{A}$ is a discrete state transition matrix, $\bar{B}$ is a discrete control input matrix, $\bar{C}$ is a discrete observation matrix, and $\bar{D}$ is a discrete feedforward matrix. ${h}_{k}$ is the current state variable, ${h}_{k-1}$ represents the hidden state of the system at the discrete time point $k$-1, containing historical information up to that time, and $k$ is the discrete time step index.
Traditional SSMs[12, 13, 2]employ static predefined parameters for matrices $\bar{A}$ , $\bar{B}$ , and $\bar{C}$ , resulting in rigid information propagation pathways. In contrast, Mamba’s SSM[11] layer implements a selective mechanism that dynamically generates timestep-specific parameters $\Delta$ conditioned on input data. This enables adaptive modulation of the discretization processes for B and C matrices, thereby establishing context-sensitive information propagation paths. Generate parameters $\Delta$, $\bar{B}$, $\bar{C}$ in real time from the input sequence through linear transformation:
${{B}_{k}}=\text{Linea}{{\text{r}}_{B}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (2)
${{C}_{k}}=\text{Linea}{{\text{r}}_{C}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (3)
${{\Delta}_{k}}=\text{softplus}(\text{Linea}{{\text{r}}_{\Delta}}({{x}_{k}}))\in{{\mathbb{R}}^{+}}$ (4)
Where: Linear represents linear projection, and $\Delta$ is the step length at which ZOH discretizes the continuous equation.
Dynamically adjust the discretization process according to the input:
${{{\bar{A}}}_{k}}=\exp({{\Delta}_{k}}{{A}_{k}})$ (5)
${{{\bar{B}}}_{k}}={{({{\Delta}_{k}}{{A}_{k}})}^{-1}}(\exp({{\Delta}_{k}}{{A}_{k}})-I){{\Delta}_{k}}{{B}_{k}}$ (6)
${{{\bar{C}}}_{k}}={{C}_{k}}$ (7)
Plugging parameter $\bar{A}$, $\bar{B}$, $\bar{C}$ into Equation 1, the selective SSM still follows the LTI property in the continuous time domain, thus allowing efficient GPU-accelerated convolution computing of the discretized state space equation:
$\bar{K}=(\overline{{{C}_{k}}{{B}_{k}}},{{\overline{{{C}_{k}}{{A}_{k}}B}}_{k}},...,\overline{{{C}_{k}}A}_{k}^{L-1}{{\overline{B}}_{k}})$ (8)
${{y}_{k}}={{x}_{k}}*\bar{K}$ (9)
Where: $\bar{K}$ is the global convolution kernel, $L$ is the length of the convolution kernel.
In summary, the selective SSM’s continuous-time equation is formally consistent with traditional SSM, but content-aware modeling is realized through the dynamic parameterization of the discretization process. This design paradigm transcends LTI limitations while inheriting the theory of continuous systems, enabling Mamba to simultaneously attain computational efficiency comparable to RNNs (O(n) complexity) and expressive power equivalent to Transformer architectures.
PointMamba[21] pioneers the introduction of Mamba architecture into point cloud processing, which significantly improves efficiency while maintaining high accuracy, demonstrating the potential of Mamba in 3D vision tasks.
3 Hybrid-Emba3D
@@ 3.1 Overview
This work establishes a hybrid framework that effectively enhances the representation performance of the Mamba. It integrates the Local geometric pooling (LGP) with geometric feature coupling and a bidirectional SSM enhanced by the collaborative feature enhancer (CoFE-BiSSM). The architecture of the model is shown in Figure 3.
Figure 3: The architecture of Hybrid-Emba3D. The input point cloud is sampled into local patches via FPS and KNN, and the coordinates and features are aggregated into embedding vectors through lightweight PointNet. After incorporating positional encoding and classification tokens, input the Hybrid-Emba3D encoder. The encoder core is composed of alternately stacked CoFE-BiSSM and LGP, supplemented by normalization and residual connections. Task-specific heads finally decode the transformed features for the classification of objects.
Given an input point cloud $P\in{{\mathbb{R}}^{N\times 3}}$ with $N$ points, the framework first initializes $L$ centroid points via Farthest Point Sampling (FPS) to form a set ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$. For each centroid ${{P}_{\text{center}}}$, the method retrieves its K-Nearest Neighbors (KNN) from the original cloud to construct local patches $x_{p}^{i}\in{{\mathbb{R}}^{K\times 3}}$ serving as structural spatial primitives. A lightweight PointNet variant subsequently performs hierarchical feature encoding on each patch, with the resultant local descriptors simultaneously serving as input token sequences for Hybrid-Emba3D architectures.
Adhering to the ViT’s[7] work, the architecture embeds a learnable [class] token preceding $L$ local tokens for global representation aggregation. Each patch ${{x}_{p}}\in{{\mathbb{R}}^{L\times K\times 3}}$ is projected and mapped into a C-dimensional feature vector ${{W}_{patch}}({{x}_{p}})\in{{R}^{L\times C}}$ as its initial patch embeddings. The [class] token is vertically stacked with all point cloud block features, forming a sequence matrix with a shape of ($L$+1)×$C$ ($L$ patch features $+$ one [class]). After superimposing position encoding ${P}_{pos}$, the geometric structure information of the original point cloud is preserved through patch position encoding, providing an initial representation for subsequent hierarchical encoding.
${{Z}^{0}}=[{{x}_{\text{cls}}}\text{+}P_{\text{pos}}^{(0)},{{W}_{patch}}(x_{p}^{1})+P_{\text{pos}}^{(1)},\cdots,{{W}_{patch}}(x_{p}^{L})+P_{\text{pos}}^{(L)}]$ (10)
@@
The encoder employs dual-phase processing (The encoder block is shown in Figure 3 (b)): For the $i$-th layer, input features ${Z}^{i-1}$ undergo normalization before the LGP component executes geometric-aware local feature extraction to generate ${{\hat{Z}}^{i}}$. Post residual summation, secondary normalization precedes the CoFE-BiSSM component, which supplants conventional attention mechanisms, enabling dynamic feature enhancement and long-range dependency capture, ultimately outputting the current layer representation ${Z}^{i}$. The overall structure retains the residual connection. The process is as follows:
${{{\hat{Z}}}^{i}}=\text{LGP}(\text{Norm}({{Z}^{i-1}}+{{P}_{\text{pos}}}))+{{Z}^{i-1}},$ (11)
${{{Z}}}^{i}=\text{CoFE-BiSSM}(\text{Norm}({{{\hat{Z}}}^{i}}))+{{{\hat{Z}}}^{i}}$ (12)
@@
3.2 Local geometric pooling with geometry-feature coupling mechanism

### v2

# HyMamba: Mamba with Hybrid Geometry-Feature Coupling for Efficient Point Cloud Classification

@@ Abstract
Point cloud classification is one of the essential technologies for achieving intelligent perception of 3D environments by machines, its core challenge is to efficiently extract local and global features. Mamba leverages state space models (SSMs) for global point cloud modeling. Although prior Mamba-based point cloud processing methods pay attention to the limitation of its flattened sequence modeling mechanism in fusing local and global features, the critical issue of weakened local geometric relevance caused by decoupling geometric structures and features in the input patches remains not fully revealed, and both jointly limit local feature extraction. Therefore, we propose HyMamba, a geometry and feature coupled Mamba framework featuring: (1) Geometry-Feature Coupled Pooling (GFCP), which achieves physically interpretable geometric information coupling by dynamically aggregating adjacent geometric information into local features; (2) Collaborative Feature Enhancer (CoFE), which enhances sparse signal capture through cross-path feature hybridization while effectively integrating global and local contexts. We conducted extensive experiments on ModelNet40 and ScanObjectNN datasets. The results demonstrate that the proposed model achieves superior classification performance, particularly on the ModelNet40, where it elevates accuracy to 95.99% with merely 0.03M additional parameters. Furthermore, it attains 98.9% accuracy on the ModelNetFewShot dataset, validating its robust generalization capabilities under sparse samples.
@@
Point cloud classification, State space model, Local geometry, Feature hybridization
@@
[label1] organization=School of Opto-electronical Engineering, Xi’an Technological University,city=Xi’an, postcode=710021, country=China
@@
[label2] organization=Xi’an Key Laboratory of Active Photoelectric Imaging Detection Technology, Xi’an Technological University, city=Xi’an, postcode=710021, country=China
@@ 1 Introduction
Point clouds directly represent the geometric features of objects through 3D coordinates, compensating for the lack of spatial information in 2D data and providing basic data support for perception tasks such as 3D object recognition and scene understanding [1]. However, the inherently disordered nature of point cloud data necessitates permutation invariant processing, and effective extraction of local geometric features is constrained by nonuniform distribution; these intrinsic characteristics present challenges for improving the accuracy and generalization capability of 3D point cloud classification models.
Early approaches discretize unordered 3D point clouds into 3D voxel grids and employ 3D convolutional networks to capture spatial features. Subsequently, PointNet [2] innovatively builds a processing framework with permutation invariance, laying the foundation for the direct representation of the disordered features of point clouds. Follow-up research further captures local geometric features by developing point cloud convolution operations [3, 4, 5]. Recent studies have introduced Transformer [6] architectures with global modeling capabilities, which utilize positional encoding and multi-head attention mechanisms to comprehensively model global dependencies within point clouds [7]. Although subsequent research has enhanced feature representation through improved local region modeling and hierarchical architecture designs, they remain constrained by the inherent quadratic computational complexity of the Transformer framework [8].
@@
To circumvent the huge computational burden generated when establishing long-range dependencies, the state space model (SSM), as a novel approach for long-sequence modeling, has been introduced into point cloud learning. It realizes efficient long sequence modeling with the mathematical characteristics of continuous state representation and linear time-invariant systems. By introducing a selective mechanism with input dependency, the Mamba model [9] innovatively constructs a state space architecture with conditional computing characteristics. It combines the timing dependence modeling capabilities of RNNs [10] with the parallel computing advantages of the Transformer to achieve linear time complexity of long-distance dependence modeling.
Pointmamba [11] pioneers the application of Mamba models to point cloud processing but neglects local context modeling. The irregular topological characteristics of point cloud data make it difficult for the global sequence modeling method of SSM to effectively capture geometric correlations, thereby overlooking local neighborhood relationships and structural details. In contrast, Mamba3D [12] catches local features of point clouds through Local Norm Pooling and bidirectional SSM (BiSSM). However, recent work E-Mamba [13] adopts local geometry pooling primarily for token reordering, which fails to establish explicit geometric-feature dependencies, limited ability to perceive local structures. More critically, the existing Mamba-based point cloud processing methods primarily focus on refining the input scanning strategy of SSMs, that is, reordering or diversifying scanning inputs. Although some methods incorporate geometric information, they serve merely to facilitate reordering without directly modeling the relationships between geometric structures and features[14, 15, 16], this may make the model more inclined to learn patterns based on fixed scan order rather than true spatial structure. In addition, SSM’s global sequential modeling struggles to directly extract hierarchical features, inevitably leading to deficient local refinement during feature extraction processes. To address these limitations, this paper proposes a Geometry-Aware and Cross-Path Feature Hybrid Enhanced Mamba Model (HyMamba). Unlike E-Mamba’s pooling strategy, our Geometric-Feature Coupled Pooling (GFCP) injects geometric priors into feature representation without extra parameters, fundamentally differentiating it from existing local pooling methods. Specifically, HyMamba realizes: (1) Lightweight local spatial enhancement via geometric-aware feature coupling; (2) Cross-path feature fusion through our Collaborative feature enhancer.
based on this unique design, our framework achieves state-of-the-art(SOTA) accuracy of 95.99% on ModelNet40, significantly outperforming existing methods and establishing new performance benchmarks as shown in Figure 1.
Figure 1: Comprehensive comparison of HyMamba with other SOTA models. (All results are based on Self-supervised Pre-training.)
@@
    We proposed an efficient point cloud classification model. This model has a strong feature learning ability, achieves extremely high classification performance.
@@
    We designed the Geometric-Feature Coupled Pooling that dynamically fuses neighbor point geometry with center features via Gaussian spatial weights, enhancing the representational ability of local geometry.
@@
    We proposed the Collaborative Feature Enhancer that addresses the absence of hierarchical features in global SSM modeling through dual-path hybrid enhancement of channel and spatial features.
  • 4.
    The experiment showed that a new record accuracy of 95.99% was achieved on ModelNet40 with 3.95G loating Point Operations (FLOPs), demonstrating excellent efficiency-accuracy balance.
@@ 2.1 Deep learning for point cloud data
The efficacy of point cloud classification hinges upon synergistic modeling of local and global contexts. Early approaches such as VoxNet [17] and OctNet [18] leverage auxiliary voxel data structures and 3D convolutions to address spatial disorder in point clouds. Nevertheless, this reliance on intermediate representations inevitably induces substantial feature degradation. To address this limitation, Qi et al.’s PointNet pioneered direct point set processing via shared multi-layer perceptrons (MLP) and symmetric functions. Building upon this foundation, PointNet++ [19] enhances local detail extraction by introducing farthest point sampling (FPS) to establish the hierarchical structure. Graph convolutional approaches further enhance local geometric feature extraction by establishing connections between neighboring points through directed graphs. For example, DGCNN [20] fuses edge features into the graph convolutional operations, consequently enhancing classification accuracy. Methodologies including Volume-based 3D convolutions [21, 22, 23], point-based direct convolutions [24, 25, 26, 27], and graph convolutional networks [20, 28, 29, 30] fundamentally rely on local feature aggregation. These methodologies fundamentally suffer from the mismatch between fixed receptive fields and irregular point distributions, which makes it difficult to model the long-distance dependency in point cloud data and limits its performance improvement in complex scenes.
@@
Owing to its excellent global modeling capabilities, Vision Transformer [31] (ViT) has become one of the mainstream architectures in point cloud analysis. Based on relative position coding, its self-attention mechanism can adaptively establish geometric correlations between point cloud elements and significantly improve the scene semantic understanding ability through global dependency modeling [32, 33, 34, 35, 36]. The conflict exists between the sparse distribution characteristics of point cloud data and the dense connectivity assumption underlying standard attention mechanisms, leading to redundant attention weight allocations. Researchers have developed sparse attention strategies to improve this. Such as DSVT [37] with hierarchical feature aggregation demonstrates optimized computation-accuracy balance through adaptive sparse attention patterns; Point-BERT [33] self-supervised pretraining via masked point modeling, effectively reducing annotation dependency; PatchFormer [38] significantly reduces computational complexity through geometrically consistent patch partitioning and streamlined patch-level attention mechanisms.
In this context, Mamba demonstrates potential for point cloud processing, achieving linear complexity in long-sequence modeling while possessing global learning ability.
2.2 Selective State Space Model
@@
SSM, a classic method in control theory, achieves system modeling by constructing the state space. As control theory evolved, the limitations of traditional input-output models have gradually emerged. SSM establishes a system state space based on input signal characteristic variables, uses state variables to characterize the dynamic evolution of the system fully, and abstracts the system into a mathematical model containing state and output equations.
Discretize the continuous-time linear time-invariant (LTI) state space model using the zero-order hold (ZOH) method. The discrete state and output equations of the discretized SSM are presented in Equation (1), with the corresponding architecture illustrated in Figure 2.
[…]
Where: $\bar{A}$, $\bar{B}$, $\bar{C}$, and $\bar{D}$ are the discrete state transition matrix, control input matrix, observation matrix, and feedforward matrix. Specifically, $\bar{A}={\exp{(\Delta A)}}$, $\bar{B}={{(\Delta A)}^{-1}}(\bar{A}-I)\Delta B$, $\ Delta$is the discrete sampling interval of ZOH. ${h}_{k}$ is the current state variable, ${h}_{k-1}$ represents the hidden state of the system at the discrete time point $k-1$, containing historical information up to that time, and $k$ is the discrete time step index.
@@
Traditional SSMs [39, 40, 41]employ static predefined parameters for matrices $\bar{A}$ , $\bar{B}$ , and $\bar{C}$ , resulting in rigid information propagation pathways. In contrast, Mamba’s SSM [9] layer implements a dynamic mechanism that dynamically adaptive modulation of ${B}_{k}$, ${C}_{k}$ and ${\Delta}_{k}$ depending input data:
${{B}_{k}}={Linear}_{B}({{x}_{k}})$ (2)
${{C}_{k}}={Linear}_{C}({{x}_{k}})$ (3)
${{\Delta}_{k}}={softplus}({Linear_{\Delta}}({{x}_{k}}))$ (4)
Where: $Linear$ represents linear projection.
Then discretize the dynamic parameters, at this time $\bar{A}_{k}={\exp({{\Delta}_{k}A})}$, ${{{\bar{B}}}_{k}}={{({{\Delta}_{k}}{{A}})}^{-1}}(\bar{A}_{k}-I){{\Delta}_{k}}{{B}_{k}}$, ${{{\bar{C}}}_{k}}={{C}_{k}}$. Plugging parameter $\bar{A}_{k}$, $\bar{B}_{k}$, $\bar{C}_{k}$ into equation (1), Selective SSMs effectively transform SSMs into time-varying systems, thereby establishing context-sensitive information propagation paths. However, this adaptation renders parallelization through convolution not viable. Consequently, Mamba introduces a parallel scan algorithm to achieve efficient parallel computing:
${{y}_{k}}$ $={\bar{C}_{k}}{{h}_{k}}={\bar{C}_{k}}Scan({{{\bar{A}}}_{k}},{{{\bar{B}}}_{k}}{{x}_{k}})$ (5)
$={\bar{C}_{k}}\left(\left(\prod\nolimits_{i=1}^{k}{\left[\begin{matrix}{{{\bar{A}}}_{i}}&{{{\bar{B}}}_{i}}{{x}_{i}}\\ 0&1\\ \end{matrix}\right]}\right)\left[\begin{matrix}0\\ 1\\ \end{matrix}\right]\right)$
Overall, the selective SSM’s continuous-time equation is formally consistent with traditional SSM, but content-aware modeling is realized through the parameter dynamization in the discretization process. This design paradigm transcends LTI limitations, enabling Mamba to simultaneously attain computational efficiency comparable to RNNs (O(n) complexity) and global expressive power equivalent to Transformer architectures in long sequence input.
3 HyMamba
@@ 3.1 Overview
This work establishes a hybrid framework that effectively enhances the representation performance of the Mamba. It integrates the Geometry-Feature Coupled Pooling (GFCP) and a bidirectional SSM enhanced by the collaborative feature enhancer (CoFE-BiSSM). The architecture of the model is shown in Figure 3.
Figure 3: The architecture of HyMamba. The input point cloud is sampled into local patches via FPS and KNN, and the coordinates and features are aggregated into embedding vectors through lightweight PointNet. After incorporating positional encoding and classification tokens, input the HyMamba encoder. The encoder core is composed of alternately stacked CoFE-BiSSM and GFCP, supplemented by normalization and residual connections. Task-specific heads finally decode the transformed features for the classification of objects.
Given an input point cloud $P\in{{\mathbb{R}}^{N\times 3}}$ with $N$ points, the framework first initializes $G$ centroid points via Farthest Point Sampling (FPS) to form a set ${{P}_{{center}}}\in{{\mathbb{R}}^{G\times 3}}$. For each centroid ${{P}_{{center}}}$, the method retrieves its K-Nearest Neighbors (KNN) from the original cloud to construct local patches $x_{p}^{i}\in{{\mathbb{R}}^{K\times 3}}$ serving as structural spatial primitives. A lightweight PointNet variant subsequently performs hierarchical feature encoding on each patch, with the resultant local descriptors simultaneously serving as input token sequences for HyMamba architectures.
Adhering to the ViT’s [31] work, the architecture embeds a learnable [class] token preceding $L$ local tokens for global representation aggregation. Each patch ${{x}_{p}}\in{{\mathbb{R}}^{G\times K\times 3}}$ is projected and mapped into a C-dimensional feature vector ${{W}_{patch}}({{x}_{p}})\in{{R}^{G\times C}}$ as its initial patch embeddings. The [class] token is vertically stacked with all point cloud block features, forming a sequence matrix with a shape of ($G$+1)×$C$ ($G$ patch features $+$ one [class]). After superimposing position encoding ${P}_{pos}$, the geometric structure information of the original point cloud is preserved through patch position encoding, providing an initial representation for subsequent hierarchical encoding.
${{Z}^{0}}=[{{x}_{{cls}}}{,}P_{{pos}}^{(0)};{{W}_{patch}}(x_{p}^{1}),P_{{pos}}^{(1)};\cdots;{{W}_{patch}}(x_{p}^{G}),P_{{pos}}^{(G)}]$ (6)
@@
The encoder employs dual-phase processing (The encoder block is shown in Figure 3 (b)): For the $i$-th layer, input features ${Z}^{i-1}$ undergo normalization before the GFCP component executes geometric-aware local feature extraction to generate ${{\hat{Z}}^{i}}$. Post residual summation, secondary normalization precedes the CoFE-BiSSM component, which supplants conventional attention mechanisms, enabling dynamic feature enhancement and long-range dependency capture, ultimately outputting the current layer representation ${Z}^{i}$. The overall structure retains the residual connection. The process is as follows:
${{{\hat{Z}}}^{i}}={GFCP}({Norm}({{Z}^{i-1}}+{{P}_{{pos}}}))+{{Z}^{i-1}},$ (7)
${{{Z}}}^{i}={CoFE-BiSSM}({Norm}({{{\hat{Z}}}^{i}}))+{{{\hat{Z}}}^{i}}$ (8)
@@
3.2 Geometry-Feature Coupled Pooling